OpenStudy (anonymous):

Help Please - Log-normal distribution The length of time (in seconds) that a user views a page on a web site before moving to another page is lognormal random variable with parameter θ = 0.5 and ω^2 = 1. 1) What is the probability that a page is viewed for more than seconds? 2) What is the length of time that 50% of users view the page?

OpenStudy (anonymous):

Hi, are you still there?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

first of all it looks like there are some numbers missing from the problem... in question 1

OpenStudy (anonymous):

nope, that's what is given

OpenStudy (anonymous):

it says "what is the probability that a page is viewed for more than seconds". how many seconds?

OpenStudy (anonymous):

oh sorry, it's 10

OpenStudy (anonymous):

ok, do you have a formula for the CDF of the lognormal distribution in your book?

OpenStudy (anonymous):

that's what i need.. can't seem to find which formula to use

OpenStudy (anonymous):

or a table of lognormal values?

OpenStudy (anonymous):

i have a Zp table

OpenStudy (anonymous):

hmm, well, I may not be able to give you a satisfying answer. the CDF of the LogNormal distribution with the parameters specified, evaluated at 10, is the probability that the time will be less than or equal to 10 seconds. So the complement of that quantity (1 - <that>) is the answer. I usually compute this by just typing it into a calculator - I don't know how your teacher expects you to calculate it by hand. but my calculator says CDF of LogNormal[0.5, 1] evaluated at 10 is 0.964, which means that the probability they're viewing for more than 10 is 0.357

OpenStudy (anonymous):

sorrry I mean 0.0357, i.e. 3.6%

OpenStudy (anonymous):

the second half of the quesiton, are you sure you typed it exactly right?

OpenStudy (anonymous):

You wrote "What is the length of time that 50% of users view the page?" - is that exactly how the question is worded?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

well that's not a very well worded question but they might be asking for the median time. then you just look up the formula for the median of the lognormal distribution which is e^u, where you said u=0.5, so it's e^0.5

OpenStudy (anonymous):

yea that's what i did E(x)

OpenStudy (anonymous):

e(x)?

OpenStudy (anonymous):

it's e^0.5, or sqrt(e)

OpenStudy (anonymous):

nvm.. that's nother question -.-

OpenStudy (anonymous):

sqrt(e) is about 1.65

OpenStudy (anonymous):

\[P(Z> \frac{\ln(a)-.5}{1}) = .5\]

OpenStudy (anonymous):

what's that

OpenStudy (anonymous):

wich came down to a = 1.649

OpenStudy (anonymous):

the length of time that 50% of users view the page

OpenStudy (anonymous):

there is no single time that 50% of users view the page

OpenStudy (anonymous):

they might be asking "what is the time such that 50% of users view the page for no longer than this amount of time"

OpenStudy (anonymous):

but in a continuous distribution like the lognormal there is no one time that you can say 50% of users viewed the page for

OpenStudy (anonymous):

hmm.. i see, the question wasn't well written i guess....

OpenStudy (anonymous):

sorry i couldn't be more help

OpenStudy (anonymous):

it was a big help. Thanks

OpenStudy (anonymous):

sure